OpenStudy (anonymous):

4m + 9 + 5m – 12 = 42

OpenStudy (jiteshmeghwal9):

combine the like terms 4m+5m-12+9=42 9m-3=42 nw just add 3 both sides & then divide 9 both sides.

OpenStudy (anonymous):

m=5

OpenStudy (anonymous):

4m + 9 + 5m - 12 = 42 (combine) 9m - 3 = 42 (add 3 to both sides) 9m = 45 (divide both sides by 9) m = 5

OpenStudy (anonymous):

thank you all soooooo much